OpenStudy (anonymous):

Express the complex number in trigonometric form. 2i Answer choices: (a)2(cos 180° + i sin 180°) (b)2(cos 0° + i sin 0°) (c)2(cos 270° + i sin 270°) (d)2(cos 90° + i sin 90°)

OpenStudy (anonymous):

2i=2(0+1i)=2(cos \(\theta+i\sin \theta\)) now you can find \( \theta\)

OpenStudy (anonymous):

\[\cos \theta=0,\theta=90,270\] \[\sin \theta=1,\theta=90\] \[\theta=?\]